OpenStudy (anonymous):

what is the solution of the following system -3x-2y=-12, 9x+6y=-9 a. (2,1) b. no solutions c.(-2,-1) d. infinitely many solutions

OpenStudy (anonymous):

substitute

OpenStudy (anonymous):

wait r u the one i helped 3 minutes ago? lol

OpenStudy (anonymous):

haha yea lol

OpenStudy (anonymous):

cool same thing here, try it yourself first

OpenStudy (anonymous):

9x+6y=-9= -1

OpenStudy (anonymous):

remember the whole point of substituting is to make an equation with only 1 variable, which we can easily solve. we cannot solve a function with more than 1 variable

OpenStudy (anonymous):

so first, choose a function, then isolate a variable, x or y, then put it in terms of the other variable, then substitute it into the other equation.

OpenStudy (anonymous):

if u choose to isolate x, u shud have an equation like x = something, u then have to put it in terms of y, meaning the "something" has only y in it, nothing else

OpenStudy (anonymous):

so say u choose -3x-2y=-12 to isolate x (or y if u want), u get x = 4 - 2/3 y, following the format i mentioned above "x = something, u then have to put it in terms of y, meaning the "something" has only y in it,"

OpenStudy (anonymous):

then plug x = 4 - 2/3 y into the second equation, then solve for y because now x is eliminated (substituted by y).

OpenStudy (anonymous):

11calcBC...do you think this problem would be more easily solved with elimination. Because if you use substitution, you will be dealing with fractions.

OpenStudy (anonymous):

yes that would be easier actually, im just trying to make clear to her the concept of substitution lol

OpenStudy (anonymous):

I understand..and you are doing a good job...thumbs up :)

OpenStudy (anonymous):

lol thanks mate!

OpenStudy (anonymous):

9x4+6x-2/3??

OpenStudy (anonymous):

show me ur equation first

OpenStudy (anonymous):

36+4???

OpenStudy (anonymous):

u know how to use elimintation?

OpenStudy (anonymous):

what is that lol?

OpenStudy (anonymous):

thats when make u directly add or subtract functions to eliminate a variable

OpenStudy (anonymous):

that would leave only 1 variable behind, which is what we always want

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

b. no solutions

OpenStudy (anonymous):

u sure

OpenStudy (anonymous):

If you use elimination\[-3x -2y = -12\]\[9x + 6y = -9\]Multiply top by 3 that gives you\[-9x - 6y = -36\]However if you add the top and bottom equations than the variables will cancel out. You will end up with: \[nothing = -45\]Therefore you have no solutions.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem.